OpenStudy (shaik0124):

A man covered a certain distance at some speed. Had he moved 3 kmph faster, he would have taken 40 minutes less. If he had moved 2 kmph slower, he would have taken 40 minutes more. The distance (in km) is: A. 35 B. 36 C. 37 D. 40

OpenStudy (anonymous):

change words to equations: s=(v+3)*(t-40).--->1 s = (v-2)*(40+t).--->2 s=v*t.---->3 then expand them out and get 1-2 . and 2-3. thus 1-2 = 2-3.

OpenStudy (shaik0124):

thank u

OpenStudy (anonymous):

noo i forget to convert them to hours t+2/3 t-2/3

OpenStudy (anonymous):

you will get from 1-2 v=(15/4)t-0.5 and from 1-3 v=3t+2. thus t = 10/3 and v = 12. so s = 40 km.

OpenStudy (shaik0124):

t+2/3 ??

OpenStudy (anonymous):

yea \[1 hr = 60 \min\] 1 ---> 60 x ___> 40 thus x = 40/60 =2/3 hr.